xx WaveToyExtra

WaveToyExtra

Cactus Maintainers <cactusmaint@cactuscode.org>

Date

Abstract

Advanced features for WaveToy

1 Introduction

This thorn contains enhancements to the standard WaveToy thorns for modelling 3D scalar fields. WaveToy is primarily intended as a example thorn and for that reason is deliberately kept as simple as practical, with more advanced features included here instead. Currently this thorn shows how more flexible boundaries can be used with WaveToy or other thorns.

2 Boundary Conditions

This thorn provides the infrastructure to use any registered external boundary condition with WaveToy, with any choice of options.

3 Parameters

11 customxbound 10 Scope: private01  STRING

11 Description: Custom boundary condition to use

11 Range 10 01 Default: (none)
11 .*  10 Any valid boundary condition

11

11 customxoptions10 Scope: private01  STRING

11 Description: Table string for custom boundary condition

11 Range 10 01 Default: (none)
11 .*  10 Any valid table string

11

11 bound  10 Scope: shared from WAVETOY01  KEYWORD

11 Extends ranges:

11 custom  10 custom boundary condition

11

4 Interfaces

General Implements: wavetoyextra

Inherits: wavetoy 5 Schedule This section lists all the variables which are assigned storage by thorn CactusWave/WaveToyExtra. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxEVOL x wavetoyextraxboundary x additional boundary conditions for 3d wave equation x

11 x 11 After:  11 wavetoyxevolution
11 x 11 Before: 11 wavetoyxboundaries
11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTRESTRICT x wavetoyextraxboundary x additional boundary conditions for 3d wave equation x

11 x 11 Before: 11 wavetoyxboundaries
11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxTERMINATE x wavetoycxterminate x tidy up x

11 x 11 Language: 11 c
11 x 11 Type:  11 function